OpenStudy (anonymous):

The sum of the ages of a brother and sister is 26. If four times the brother's age is subtracted from three times the sister's age, the difference is 29. What are the ages of the brother and sister? I got this far: 29 = 3s-4b

OpenStudy (anonymous):

and the sum of ages is 26 -that's your second equation

OpenStudy (gw2011):

You actually have two equations to solve. Let B=Brother's age Let S=Sister's age 3S-4B=29 S+B=26 Eliminate one of the unknowns - in this case I chose B. To do this multiply the second equation by 4 and you will get the following equations to solve: 3S-4B=29 4S+4B=104 (add these two equations): 7S=133 S=133/7 S=19 (substitute this value into one of the equations - I chose the second one): S+B=26 19+B=26 B=26-19 B=7 Therefore, the sister's age is 19 and the brother's age is 7. You can verify these ages by substituting them into both equations.

OpenStudy (anonymous):

I forgot what substitution was so this was a refresher on my memory. Thank you!

OpenStudy (gw2011):

You're welcome

OpenStudy (anonymous):

@gw2011 http://openstudy.com/code-of-conduct